Find the slope of the line that passes through(-2,9.5) and (4,5).
Answer by nerdybill(7384)

Find the slope of the line that passes through(-2,9.5) and (4,5).
.
m (slope) = (y2 - y1)/(x2 - x1)
.
Plugging in our values into:
m = (y2 - y1)/(x2 - x1)
m = (9.5 - 5)/(-2.9 - 4)
m = (9.5 - 5)/(-2 - 4)
m = (4.5)/(-6)
m = - 4.5/6
m = - 1.5/2
m = - (3/2)/2
m = - 3/4